# Divergence Minimization Preference Optimization for Diffusion Model Alignment

### v1

@@ Abstract
Diffusion models have achieved remarkable success in generating realistic and versatile images from text prompts. Inspired by the recent advancements of language models, there is an increasing interest in further improving the models by aligning with human preferences. However, we investigate alignment from a divergence minimization perspective and reveal that existing preference optimization methods are typically trapped in suboptimal mean-seeking optimization. In this paper, we introduce Divergence Minimization Preference Optimization (DMPO), a novel and principled method for aligning diffusion models by minimizing reverse KL divergence, which asymptotically enjoys the same optimization direction as original RL. We provide rigorous analysis to justify the effectiveness of DMPO, and conduct comprehensive experiments to validate its empirical strength across both human evaluations and automatic metrics. Our extensive results show that diffusion models fine-tuned with DMPO can consistently outperform or match existing techniques, specifically outperforms all existing diffusion alignment baselines by at least 64.6% in PickScore across all evaluation datasets, demonstrating the method’s superiority in aligning generative behavior with desired outputs. Overall, DMPO unlocks a robust and elegant pathway for preference alignment, bridging principled theory with practical performance in diffusion models.
@@ 1 Introduction
Diffusion models ddpm ; sohl2015deep ; song2019generative ; song2021scorebased have emerged as a leading approach for text-to-image (T2I) generation ramesh2022hierarchical ; pernias2024wrstchen ; zsimg , offering a scalable and robust framework for synthesizing images from natural language descriptions. These models typically adopt a single-stage training approach, learning the data distribution from large-scale web-crawled text-image pairs sd15 ; sdxl ; dit . In contrast, Large Language Models (LLMs) have demonstrated remarkable success using a two-stage training paradigm (gpt4, ; llama3, ). After pretraining on vast and noisy web data, LLMs undergo a crucial second phase of fine-tuning on smaller but more specialized datasets. This two-stage approach allows the models to first develop broad capabilities and then align with user preferences to fulfill diverse human needs (qwen2025, ; llama3, ). The fine-tuning phase enhances model responsiveness to human expectations without significantly diminishing the broad capabilities acquired during pretraining. These alignment techniques have provided tremendous inspiration for the text-to-image domain, where leveraging human preference data for a given prompt holds great promise of unlocking diffusion models’ ability to align with human expectations. However, diffusion models and LLMs operate under fundamentally different mechanisms—LLMs rely on autoregressive factorization while diffusion models are built with a chain of Markov transitions. As a result, porting these two-stage alignment methods to diffusion architectures remains a significant challenge.
@@
In recent years, various RL-based methods have been developed for diffusion model alignment (traindiffrl, ; emu3, ; diffusiondpo, ; kto, ; directft, ; backrl, ). Early-stage works focused on fine-tuning diffusion models through Reinforcement Learning from Human Feedback (RLHF) (rlhf, ) after large-scale pretraining (traindiffrl, ; directft, ; emu3, ). These approaches typically first fit a reward model on human preference data, and then optimize the diffusion model to generate images that receive high reward scores while avoiding excessive deviation from the original model. However, building reliable reward models for diverse tasks presents challenges, often requiring substantial datasets and considerable training resources (wang2024helpsteer, ; hpsv2, ; imagereward, ). To address this issue, recent research has focused on alignment techniques without reliance on reward models. Diffusion-DPO (diffusiondpo, ) was the first to avoid the reward model by extending the formulation of direct preference optimization (DPO) to diffusion models. However, our analysis reveals that this method corresponds to minimizing the variational bound of the forward Kullback-Leibler (KL) divergence between the target and model distributions, which is known to encourage compromised mean-seeking behavior (minka2005divergence, ; exo, ). As a result, the resulting model often tends to cover diverse human intent but fails to capture exact modes of human preference, resulting in blurry or diluted generations. Followup works (kto, ; dspo, ) further propose learning with binary feedback instead of preference comparison, or fine-tuning the diffusion models from a score matching perspective. Nevertheless, none of these models challenge DPO’s underlying formulation and thus the alignment performance degrades to a weak alignment method similar to supervised fine-tuning (SFT). As a result, developing principled and exact alignment methods for diffusion models remains an open problem. In this paper, we first provide the formal analysis that the Diffusion-DPO objective is equivalent to optimizing the variational upper bound of KL divergence between target and model distributions. Based on our observation, we revisit the diffusion alignment framework from the new distribution matching perspective and argue that, compared to minimizing forward KL, optimizing reverse KL divergence provides a more mode-seeking objective. Such an objective enables more precise optimization towards the major mode in target distribution, thereby more accurately capturing the true structure of human preferences. Based on this insight, we propose Divergence Minimization Preference Optimization (DMPO), a method that fine-tunes diffusion models by minimizing the reverse KL between the model distribution and the theoretical optimal distribution. Though sharing the same optimal solution as Diffusion-DPO, in the practical scenario with limited network expressivity, DMPO enables higher human preferences alignment behavior by pushing the probability mass to the main characteristics of the target distribution. Furthermore, we theoretically show that under mild assumptions, DMPO optimizes the policy distribution in the same direction as the original RL objective which provides justification for the rationality and effectiveness of our proposed algorithm.
@@ 2.1 Diffusion Models
Suppose real data sample ${\mathbf{x}}_{0}$ follows data distribution $q({\mathbf{x}}_{0})$, denoising diffusion models (ddpm, ) are generative models which have a discrete-time reverse process with a Markov structure $p_{\theta}({\mathbf{x}}_{0:T})=\prod_{t=1}^{T}p_{\theta}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ where
@@
Training the models is performed by minimizing the associated evidence lower bound (ELBO) (vaediffusion, ; Song2021MaximumLT, ):
$L_{\text{Diff}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t,{\mathbf{x}}_{t}}\left[\omega(\lambda_{t})\|\bm{\epsilon}-\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\|_{2}^{2}\right],$ (2)
with $\bm{\epsilon}\!\sim\!\mathcal{N}(0,\bm{I})$, $t\!\sim\!\mathcal{U}(0,T)$,
${\mathbf{x}}_{t}\sim q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t};\alpha_{t}{\mathbf{x}}_{0},\sigma^{2}_{t}\bm{I}).$ $\lambda_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$ is a signal-to-noise ratio (vaediffusion, ), and $\omega(\lambda_{t})$ is a pre-specified weighting function (typically chosen to be constant (ddpm, ; song2019generative, )).
@@ 2.2 Reinforcement Learning from Human Feedback (RLHF)
RLHF is a foundational paradigm for aligning large-scale generative models with human preferences. The standard RLHF pipeline consists of three stages: (1) supervised fine-tuning (SFT) on curated instruction data, (2) learning a reward model from human preference comparisons, and (3) reinforcement learning to fine-tune the policy using the learned reward. In the RL phase, a parameterized policy $p_{\theta}({\mathbf{x}}_{0}|c)$ (input condition $c\sim\mathcal{D}_{c}$) is optimized to maximize expected reward under the guidance of a fixed reward function $r(c,{\mathbf{x}}_{0})$, typically trained using a Bradley–Terry model btmodel on human-labeled preference pairs. To prevent the policy from drifting too far from a reference distribution $p_{\text{ref}}({\mathbf{x}}_{0}|c)$ (often obtained from the SFT stage) the objective includes a KL-regularization term:
@@
where $\beta>0$ is a hyperparameter controlling regularization. Analytically, the optimal policy under this objective corresponds to a Boltzmann-rational distribution over the reference model optreward :
@@
and substituting into the preference-based objective Equation 3 yields the DPO loss:
@@ 3.1 Revisiting Diffusion-DPO
Diffusion-DPO is the first method to apply DPO to diffusion models. The approach proposes to minimize Equation 3 of diffusion models by replacing the KL-divergence term with its upper bound $\mathbb{D}_{\mathrm{KL}}\left[p_{\theta}(\mathbf{x}_{0:T}|c)\,\|\,p_{\mathrm{ref}}(\mathbf{x}_{0:T}|c)\right]$, which is joint KL-divergence on sampling path ${\mathbf{x}}_{0:T}$. Based on the binary DPO formulation Equation 6, given text prompt $c$, image pairs ${\mathbf{x}}^{w}_{0},{\mathbf{x}}^{\ell}_{0}$, Diffusion-DPO formulates the RLHF objective into the following objective:
@@
which allows the diffusion model to be optimized using pairwise preference data. By leveraging the convexity of the $-\log\sigma$ function and further simplifying the DPO objective across two denoising trajectories, such loss can be decomposed as per-step alignment loss. This enables the model to efficiently align with human preferences through direct optimization of the denoising process.
@@
(informal) Generalizing Diffusion-DPO from the pairwise preference setting to the multi-sample setting with preference data sampled from the reference policy $p_{\text{ref}}$, we have that the gradient of Diffusion-DPO objective Equation 7 satisfies:
@@ 3.2 Our Method
While DPO optimizes a forward KL divergence between the optimal policy and the learned policy, this formulation introduces a key limitation: it enforces a mean-seeking behavior. Specifically, minimizing the Forward KL penalizes the learned policy $p_{\theta}$ harshly whenever it assigns low probability to any output favored by the optimal policy $p^{*}$. This encourages $p_{\theta}$ to cover the full support of $p^{*}$, often resulting in diffuse generations that fail to concentrate on highly preferred samples. In contrast, minimizing the reverse KL divergence provides a more targeted form of alignment. The reverse KL is given by $\mathbb{D}_{\mathrm{KL}}\bigl{(}\hat{p}_{\theta}({\mathbf{x}}_{0:T}|c)\,\|\,\hat{p}^{*}({\mathbf{x}}_{0:T}|c)\bigr{)}$, and only penalizes the model when it places mass in regions unsupported by $p^{*}$. This drives the model toward the high-reward modes of $p^{*}$, promoting sharper and more precise alignment with human preferences.
Based on this insight, we propose Divergence Minimization Preference Optimization (DMPO), which replaces the forward KL with a reverse KL objective to align diffusion models with human preferences by minimizing a divergence between the learned policy and the latent optimal policy derived from Equation 4. According to the optimal policy $p^{*}$ in Equation 4, we can rewrite the reward function in the form of $r(c,{\mathbf{x}}_{0:T})=\log\hat{p}^{*}({\mathbf{x}}_{0:T}|c)-\log p_{\text{ref}}({\mathbf{x}}_{0:T}|c)$. Besides, we can define the relative logits $g_{\theta}(c,{\mathbf{x}}_{0:T})=\log\hat{p}_{\theta}({\mathbf{x}}_{0:T}|c)-\log p_{\text{ref}}({\mathbf{x}}_{0:T}|c)=\beta(\log p_{\theta}({\mathbf{x}}_{0:T})-\log p_{\text{ref}}({\mathbf{x}}_{0:T}))$ Putting $r(c,{\mathbf{x}}_{0:T})$ and $g_{\theta}(c,{\mathbf{x}}_{0:T})$ into $\mathbb{D}_{\mathrm{KL}}\bigl{(}\hat{p}_{\theta}({\mathbf{x}}_{0:T}|c)\,\|\,\hat{p}^{*}({\mathbf{x}}_{0:T}|c)\bigr{)}$, the divergence objective becomes:
@@
Learning with Pair-wise Preference Data. We consider the practical objective given a dataset $\mathcal{D}$ consisting of text prompts $c$ and preference-labeled output pairs $({\mathbf{x}}^{w}_{0},{\mathbf{x}}^{\ell}_{0})$, where ${\mathbf{x}}^{w}_{0}$ is preferred over ${\mathbf{x}}^{\ell}_{0}$. By setting $K=2$ in Equation 10, Equation 9 can be expressed as:”
@@
Since $p_{\theta}({\mathbf{x}}_{0:T})$ is defined over the full trajectory $\mathbf{x}_{0:T}$ is computationally intractable, and the marginal likelihood $p_{\theta}(\mathbf{x}_{0:T})$ can not be evaluated exactly, to address this, we instead aim to estimate an upper bound of the original objective by applying Jensen’s inequality. Before doing so, we examine the convexity of the function $f(u)$ which is extracted from Equation 11:
$f(u)=\sigma(u)\log\frac{\sigma(u)}{1-\alpha}+\sigma(-u)\log\frac{\sigma(-u)}{\alpha}.$ (12)
To analyze the convexity of Equation 12, we compute its second derivative $f^{\prime\prime}(u)$. Since $\alpha$ should be a very small positive constant, we find that there always exists an interval $(h_{1}(\alpha),h_{2}(\alpha))$ within which the function $f(u)$ maintains a strictly positive second derivative (e.g., when $\alpha=10^{-3}$, $h_{1}(\alpha)=-0.28$, $h_{2}(\alpha)=7.90$). In practice, the model does not choose preference data and is typically initialized near $u=0$. As the optimization progresses, the optimization gradually pushes $u$ in the positive direction, as the model increasingly focuses on modeling the winner distribution. The optimal value is reached when $u\to\log\frac{1-\alpha}{\alpha}$, which corresponds to the point where the model prediction matches the prior preference distribution. We also verify that for any $\alpha\in(0,1)$, the right endpoint $h_{2}(\alpha)$ is always greater than $\log\frac{1-\alpha}{\alpha}$. Hence, the practical optimization trajectory $u\in(0,\log\frac{1-\alpha}{\alpha})$ always lies within the convexity-preserving interval $(h_{1}(\alpha),h_{2}(\alpha))$.
We made a mild assumption that the interval $(h_{1}(\alpha),h_{2}(\alpha))$ guaranteeing the convexity of the function is both wide enough for alignment. A more detailed theoretical derivation of this interval as a function of $\alpha$ is provided in Appendix A. Under this assumption we can use the jensen inequality to move the expectation of ${\mathbf{x}}$ in Equation 11 outside:
@@
where $u_{t}(\theta)$ is defined as the following form:
@@
$=$ $-\beta T\omega(\lambda_{t})\Big{(}\|\epsilon^{w}-\epsilon_{\theta}({\mathbf{x}}_{t}^{w},t)\|_{2}^{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}({\mathbf{x}}_{t}^{w},t)\|_{2}^{2}-\left(\|\epsilon^{l}-\epsilon_{\theta}({\mathbf{x}}_{t}^{l},t)\|_{2}^{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}({\mathbf{x}}_{t}^{l},t)\|_{2}^{2}\right)\Big{)}$

### v2

@@ Abstract
Diffusion models have achieved remarkable success in generating realistic and versatile images from text prompts. Inspired by the recent advancements of language models, there is an increasing interest in further improving the models by aligning with human preferences. However, we investigate alignment from a divergence minimization perspective and reveal that existing preference optimization methods are typically trapped in suboptimal mean-seeking optimization. In this paper, we introduce Divergence Minimization Preference Optimization (DMPO), a novel and principled method for aligning diffusion models by minimizing reverse KL divergence, which asymptotically enjoys the same optimization direction as original RL. We provide rigorous analysis to justify the effectiveness of DMPO, and conduct comprehensive experiments to validate its empirical strength across both human evaluations and automatic metrics. Our extensive results show that diffusion models fine-tuned with DMPO can consistently outperform or match existing techniques, specifically consistently outperforming all baseline models across different base models and test sets, achieving the best PickScore in every case, demonstrating the method’s superiority in aligning generative behavior with desired outputs. Overall, DMPO unlocks a robust and elegant pathway for preference alignment, bridging principled theory with practical performance in diffusion models.
@@ 1 Introduction
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Song et al., 2021b) have emerged as a leading approach for text-to-image (T2I) generation (Ramesh et al., 2022; Pernias et al., 2024; Ramesh et al., 2021), offering a scalable and robust framework for synthesizing images from natural language descriptions. These models typically adopt a single-stage training approach, learning the data distribution from large-scale web-crawled text-image pairs (Rombach et al., 2022; Podell et al., 2023; Peebles and Xie, 2023). In contrast, Large Language Models (LLMs) have demonstrated remarkable success using a two-stage training paradigm (Achiam et al., 2023; Dubey et al., 2024). After pretraining on vast and noisy web data, LLMs undergo a crucial second phase of fine-tuning on smaller but more specialized datasets. This two-stage approach allows the models to first develop broad capabilities and then align with user preferences to fulfill diverse human needs (Yang et al., 2024b; Dubey et al., 2024). The fine-tuning phase enhances model responsiveness to human expectations without significantly diminishing the broad capabilities acquired during pretraining. These alignment techniques have provided tremendous inspiration for the text-to-image domain, where leveraging human preference data for a given prompt holds great promise of unlocking diffusion models’ ability to align with human expectations. However, diffusion models and LLMs operate under fundamentally different mechanisms—LLMs rely on autoregressive factorization while diffusion models are built with a chain of Markov transitions. As a result, porting these two-stage alignment methods to diffusion architectures remains a significant challenge.
In recent years, various RL-based methods have been developed for diffusion model alignment (Miao et al., 2024; Dai et al., 2023; Wallace et al., 2024; Li et al., 2024; Prabhudesai et al., 2023). Early-stage works focused on fine-tuning diffusion models through Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) after large-scale pretraining (Miao et al., 2024; Dai et al., 2023). These approaches typically first fit a reward model on human preference data, and then optimize the diffusion model to generate images that receive high reward scores while avoiding excessive deviation from the original model. However, building reliable reward models for diverse tasks presents challenges, often requiring substantial datasets and considerable training resources (Wang et al., 2024; Wu et al., 2023; Xu et al., 2023). To address this issue, recent research has focused on alignment techniques without reliance on reward models. Diffusion-DPO (Wallace et al., 2024) was the first to avoid the reward model by extending the formulation of direct preference optimization (DPO) to diffusion models.
@@
However, our analysis reveals that this method corresponds to minimizing the variational bound of the forward Kullback-Leibler (KL) divergence between the target and model distributions, which is known to encourage compromised mean-seeking behavior (Minka et al., 2005; Ji et al., 2024). As a result, the resulting model often tends to cover diverse human intent but fails to capture exact modes of human preference, resulting in blurry or diluted generations. Followup works (Li et al., 2024; Zhu et al., 2025) further propose learning with binary feedback instead of preference comparison, or fine-tuning the diffusion models from a score matching perspective. Nevertheless, none of these models challenge DPO’s underlying formulation and thus the alignment performance degrades to a weak alignment method similar to supervised fine-tuning (SFT). As a result, developing principled and exact alignment methods for diffusion models remains an open problem. In this paper, we first provide the formal analysis that the Diffusion-DPO objective is equivalent to optimizing the variational upper bound of KL divergence between target and model distributions. Based on our observation, we revisit the diffusion alignment framework from the new distribution matching perspective and argue that, compared to minimizing forward KL, optimizing reverse KL divergence provides a more mode-seeking objective. Such an objective enables more precise optimization towards the major mode in target distribution, thereby more accurately capturing the true structure of human preferences. Based on this insight, we propose Divergence Minimization Preference Optimization (DMPO), a method that fine-tunes diffusion models by minimizing the reverse KL between the model distribution and the theoretical optimal distribution. Though sharing the same optimal solution as Diffusion-DPO, in the practical scenario with limited network expressivity, DMPO enables higher human preferences alignment behavior by pushing the probability mass to the main characteristics of the target distribution. Furthermore, we theoretically show that under mild assumptions, DMPO optimizes the policy distribution in the same direction as the original RL objective which provides justification for the rationality and effectiveness of our proposed algorithm.
@@ 2.1 Diffusion Models
Suppose real data sample ${\mathbf{x}}_{0}$ follows data distribution $q({\mathbf{x}}_{0})$, denoising diffusion models (Ho et al., 2020) are generative models which have a discrete-time reverse process with a Markov structure $p_{\theta}({\mathbf{x}}_{0:T})=\prod_{t=1}^{T}p_{\theta}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ where
@@
Training the models is performed by minimizing the associated evidence lower bound (ELBO) (Kingma et al., 2021; Song et al., 2021a):
${\mathcal{L}}_{\text{Diff}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t,{\mathbf{x}}_{t}}\left[\omega(\lambda_{t})\|\bm{\epsilon}-\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\|_{2}^{2}\right],$ (2)
with $\bm{\epsilon}\!\sim\!\mathcal{N}(0,\bm{I})$, $t\!\sim\!\mathcal{U}(0,T)$, ${\mathbf{x}}_{t}\sim q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t};\alpha_{t}{\mathbf{x}}_{0},\sigma^{2}_{t}\bm{I}).$ $\lambda_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$ is a signal-to-noise ratio (Kingma et al., 2021), and $\omega(\lambda_{t})$ is a pre-specified weighting function (typically chosen to be constant (Ho et al., 2020; Song and Ermon, 2019)).
@@ 2.2 Reinforcement Learning from Human Feedback (RLHF)
RLHF is a foundational paradigm for aligning large-scale generative models with human preferences. The standard RLHF pipeline consists of three stages: (1) supervised fine-tuning (SFT) on curated instruction data, (2) learning a reward model from human preference comparisons, and (3) reinforcement learning to fine-tune the policy using the learned reward. In the RL phase, a parameterized policy $p_{\theta}({\mathbf{x}}_{0}|c)$ (input condition $c\sim\mathcal{D}_{c}$) is optimized to maximize expected reward under the guidance of a fixed reward function $r(c,{\mathbf{x}}_{0})$, typically trained using a Bradley–Terry model (Bradley and Terry, 1952) on human-labeled preference pairs. To prevent the policy from drifting too far from a reference distribution $p_{\text{ref}}({\mathbf{x}}_{0}|c)$ (often obtained from the SFT stage) the objective includes a KL-regularization term:
@@
where $\beta>0$ is a hyperparameter controlling regularization. Analytically, the optimal policy under this objective corresponds to a Boltzmann-rational distribution over the reference model (Peters and Schaal, 2007):
@@
and substituting into the preference-based objective Equation˜3 yields the DPO loss:
@@ 3.1 Revisiting Diffusion-DPO
Diffusion-DPO is the first method to apply DPO to diffusion models. The approach proposes to minimize Equation˜3 of diffusion models by replacing the KL-divergence term with its upper bound $\mathbb{D}_{\mathrm{KL}}\left[p_{\theta}(\mathbf{x}_{0:T}|c)\,\|\,p_{\mathrm{ref}}(\mathbf{x}_{0:T}|c)\right]$, which is joint KL-divergence on sampling path ${\mathbf{x}}_{0:T}$. Based on the binary DPO formulation Equation˜6, given text prompt $c$, image pairs ${\mathbf{x}}^{w}_{0},{\mathbf{x}}^{\ell}_{0}$, Diffusion-DPO formulates the RLHF objective into the following objective:
@@
which allows the diffusion model to be optimized using pairwise preference data. By leveraging the convexity of the $\log\sigma$ function and further simplifying the DPO objective across two denoising trajectories, such loss can be decomposed as per-step alignment loss. This enables the model to efficiently align with human preferences through direct optimization of the denoising process.
@@
(informal) Generalizing Diffusion-DPO from the pairwise preference setting to the multi-sample setting with preference data sampled from the reference policy $p_{\text{ref}}$, we have that the gradient of Diffusion-DPO objective Equation˜7 satisfies:
@@ 3.2 Our Method
Figure 1: Visualization of the effect of forward KL (DiffusionDPO) and reverse KL (DMPO) alignment on policy learning. We sample from MLP diffusion models trained with DPO and DMPO objectives. Green contours indicate the desirable distribution, while red contours denote the undesirable distribution. Compared to DiffusionDPO (orange samples), DMPO (blue samples) aligns more accurately with the target by concentrating on the dominant mode of the mixture Gaussian, highlighting its stronger alignment capability.
While DPO optimizes a forward KL divergence between the optimal policy and the learned policy, this formulation introduces a key limitation: it enforces a mean-seeking behavior. Specifically, minimizing the Forward KL penalizes the learned policy $p_{\theta}$ harshly whenever it assigns low probability to any output favored by the optimal policy $p^{*}$. This encourages $p_{\theta}$ to cover the full support of $p^{*}$, often resulting in diffuse generations that fail to concentrate on highly preferred samples. In contrast, minimizing the reverse KL divergence provides a more targeted form of alignment. The reverse KL is given by $\mathbb{D}_{\mathrm{KL}}\bigl(\hat{p}_{\theta}({\mathbf{x}}_{0:T}|c)\,\|\,\hat{p}^{*}({\mathbf{x}}_{0:T}|c)\bigr)$, and only penalizes the model when it places mass in regions unsupported by $p^{*}$. This drives the model toward the high-reward modes of $p^{*}$, promoting sharper and more precise alignment with human preferences. Figure˜1 provides a toy 2D illustration of this contrast, where forward KL exhibits mean-seeking behavior while reverse KL concentrates on the dominant mode. A more detailed analysis is provided in Section˜5.1.
Based on this insight, we propose Divergence Minimization Preference Optimization (DMPO), which replaces the forward KL with a reverse KL objective to align diffusion models with human preferences by minimizing a divergence between the learned policy and the latent optimal policy derived from Equation˜4. According to the optimal policy $p^{*}$ in Equation˜4, we can rewrite the reward function in the form of $r(c,{\mathbf{x}}_{0:T})=\log\hat{p}^{*}({\mathbf{x}}_{0:T}|c)-\log p_{\text{ref}}({\mathbf{x}}_{0:T}|c)$. Besides, we can define the relative logits $g_{\theta}(c,{\mathbf{x}}_{0:T})=\log\hat{p}_{\theta}({\mathbf{x}}_{0:T}|c)-\log p_{\text{ref}}({\mathbf{x}}_{0:T}|c)=\beta(\log p_{\theta}({\mathbf{x}}_{0:T})-\log p_{\text{ref}}({\mathbf{x}}_{0:T}))$ . Putting $r(c,{\mathbf{x}}_{0:T})$ and $g_{\theta}(c,{\mathbf{x}}_{0:T})$ into $\mathbb{D}_{\mathrm{KL}}\bigl(\hat{p}_{\theta}({\mathbf{x}}_{0:T}|c)\,\|\,\hat{p}^{*}({\mathbf{x}}_{0:T}|c)\bigr)$, the divergence objective becomes:
@@
Learning with Pair-wise Preference Data. We consider the practical objective given a dataset $\mathcal{D}$ consisting of text prompts $c$ and preference-labeled output pairs $({\mathbf{x}}^{w}_{0},{\mathbf{x}}^{\ell}_{0})$, where ${\mathbf{x}}^{w}_{0}$ is preferred over ${\mathbf{x}}^{\ell}_{0}$. By setting $K=2$ in Equation˜10, Equation˜9 can be expressed as:
@@
Since $u(\theta)$ is an expectation over the trajectory $\mathbf{x}_{1:T}$, it is computationally intractable. To address this, we instead aim to estimate an upper bound of the original objective by applying Jensen’s inequality. To this end, we analyze the convexity of the function $f(u)$ from Equation˜11:
$f(u)=\sigma(u)\log\tfrac{\sigma(u)}{1-\alpha}+\sigma(-u)\log\tfrac{\sigma(-u)}{\alpha}.$ (12)
By analyzing the second derivative $f^{\prime\prime}(u)$, we observe that for small $\alpha>0$, there always exists an interval $(h_{1}(\alpha),h_{2}(\alpha))$ where $f(u)$ is strictly convex. In practice, the model is initialized near $u=0$ and optimization drives $u$ positively as it emphasizes the winner distribution. The optimum occurs at $u^{\ast}=\log\tfrac{1-\alpha}{\alpha}$, and for any $\alpha\in(0,1)$, we verify that $h_{2}(\alpha)>u^{\ast}$, ensuring that the practical optimization trajectory $u\in(0,u^{\ast})$ remains within the convexity-preserving region.
This analysis relies on the mild assumption that $(h_{1}(\alpha),h_{2}(\alpha))$ is sufficiently wide for alignment. A detailed derivation of this interval as a function of $\alpha$ is provided in Appendix˜A. Under this assumption we can use the Jensen inequality to move the expectation of $\mathbf{x}$} in Equation˜11 outside:
@@
where $u_{t}(\theta)$ is further simplified into the following form:
@@
$=$ $-\beta T\omega(\lambda_{t})\Big(\big(\|\epsilon^{w}-\epsilon_{\theta}({\mathbf{x}}_{t}^{w},t)\|_{2}^{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}({\mathbf{x}}_{t}^{w},t)\|_{2}^{2}\big)-\big(\|\epsilon^{l}-\epsilon_{\theta}({\mathbf{x}}_{t}^{l},t)\|_{2}^{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}({\mathbf{x}}_{t}^{l},t)\|_{2}^{2}\big)\Big)$